# Realistic Environmental Injection Attacks on GUI Agents

### v1

@@ 3. Threat Model
Attacker’s Capabilities. Attackers can upload trigger images to a target website. Because website layouts and styles typically remain stable over short periods, attackers can reliably anticipate the overall structure and appearance of the target page. Furthermore, we assume that attackers have white-box access to the model powering the GUI agent, including knowledge of its gradients and architecture  [^1] We also consider transfer-based black-box attacks in our evaluation.. Given the increasing deployment of GUI agents and the ease with which trigger images can be distributed online (Wang et al., 2024b; Tur et al., 2025), compromising even one popular model may lead to far-reaching consequences. In practice, even if only a small fraction of users rely on the targeted model, such attacks can still result in significant security risks across the broader internet ecosystem.
4. Our Motivations
(a) Successful case
(b) Unsuccessful case
Under the threat model defined in Section 3, existing environmental injection attacks are either infeasible or exhibit low effectiveness. In this section, we analyze two major challenges faced when attacking GUI agents, motivating our proposed attack method.
Challenge (1) The position and surrounding content of the trigger are unpredictable, making successful attacks in unknown contexts difficult.
Due to the dynamic nature of internet content, attackers cannot anticipate the exact position of their uploaded trigger image or the surrounding visual content. For example, a trigger image could appear at the top-right corner alongside other products of the same category in one user’s view, while in another user’s view, the same trigger image might be centrally placed among best-selling items. Ensuring effective attacks across such diverse and dynamic visual contexts is challenging.
Achieving successful attacks in dynamic environments hinges on acquiring a realistic and diverse set of training data for the target website (Brown et al., 2017). However, manually collecting webpage samples with diverse contexts for training is often laborious and time-consuming (Kasuga and Yonetani, 2024). To overcome this significant hurdle, we introduce LLM-Driven Environment Simulation. This innovative approach leverages the powerful generative capabilities of Large Language Models to automatically synthesize highly realistic and diverse training data. By training the trigger image under these varied contexts, we effectively enhance its generalizability, thereby enabling successful attacks even in dynamic and unknown environments. Detailed descriptions can be found in Section 5.2.
Challenge (2) The trigger image constitutes only a small fraction of the overall website screenshot, making it difficult for the model to consistently focus on it.
Although our experiments show that the attack effectiveness improves substantially when LLM-Driven Environment Simulation is employed during training, the overall performance still falls short of expectations. We attribute this to the limited size of the trigger image relative to the entire webpage screenshot, typically constituting around 5%, thereby limiting the trigger image’s ability to consistently affect the agent’s decision-making process.
We hypothesize that consistently directing the model’s attention to the trigger region is crucial for overcoming this challenge (Lu et al., 2025). Empirical results obtained from evaluating trigger images trained solely with LLM-Driven Environment Simulation support this hypothesis. Specifically, in about 82% of successful attack cases, the average attention weight on the trigger image region is higher than that in other regions. In contrast, this proportion drops to roughly 61% in failed cases, a value that is close to a random distribution. Figure 1 also exemplifies this, showing significant attention concentration on the trigger region in successful attack case, while attention is dispersed evenly across the screenshot in unsuccessful case.
Motivated by these insights, we introduce Attention Black Hole, which explicitly encourages the model’s attention to converge on the trigger image region. This technique significantly enhances attack effectiveness, effectively addressing the second challenge. Further details can be found in Section 5.3.
@@ 5. Methodology
To address the above challenges, we propose Chameleon, an effective approach for attacking existing GUI agents. We first detail the overall attack procedure. Then, we provide an in-depth explanation of our LLM-Driven Environment Simulation, which enables the automatic generation of realistic and dynamic training data. Finally, we introduce Attention Black Hole, designed to encourage the model’s attention to converge on the trigger image region. The overview of Chameleon is depicted in Figure 2.
@@ 5.1. Overview
For a selected target website, such as a social media platform like RedNote  [^2] https://www.xiaohongshu.com/ or an e-commerce site like Amazon  [^3] https://www.amazon.com/, an attacker first chooses a raw trigger image $I$, such as a product image or a post cover that appears normal and benign, which will then be optimized during subsequent training.
At each training step, we first employ our proposed LLM-Driven Environment Simulation $G_{LES}$ (Section 5.2), to sample a new realistic context. This provides a webpage screenshot $s$ embedded with the perturbed trigger image, a corresponding user instruction $p_{u}$, and a pixel-level mask $mask^{shot}$ that identifies the trigger’s location within $s$. The perturbation $\delta$ within the trigger image region is then updated using Projected Gradient Descent (PGD) (Madry et al., 2017). To ensure the resulting malicious trigger image remains imperceptible to users, we constrain the perturbation $\delta$ to a predefined $\ell_{\infty}$ norm bound $\epsilon$, formally:
@@
where $\lambda$ is a hyperparameter used to balance the influence between these two loss terms.
After the training procedure described above, the attacker uploads the resulting trigger image to the target website, acting as a regular user. Once the trigger image is integrated into a public post or product listing, any GUI agent browsing the site and encountering the trigger image within a webpage will be susceptible to attack, potentially performing the intended harmful action. Algorithm 1 summarizes the overall procedure of Chameleon, covering the preprocessing, training, and deployment stages.
@@
$C\leftarrow$ Crawl large-scale multi-modal content;
@@ 5.2. LLM-Driven Environment Simulation
To address the challenge posed by the variability in trigger positions and their surrounding content, we introduce LLM-Driven Environment Simulation (LES). This approach leverages known webpage layouts and styles, employing large language models to automatically synthesize realistic and diverse training samples that capture authentic contexts. Specifically, we define a generation function $G_{LES}$, which outputs a rendered webpage screenshot, a binary mask indicating the location of the trigger image, and contextually relevant user instructions for downstream training.
First, to ensure our simulated environments are realistic, we create a high-fidelity HTML template $T$ for each target website. We leverage a flagship LLM to parse the live website, preserving its core structural and stylistic components while removing existing content. This results in a template that is visually authentic.
We then construct a large-scale, multi-modal corpus $C$ by crawling the target website for actual content, gathering an average of over 5,000 distinct items (e.g., product images, titles, prices) for each website. This process ensures our content pool is not only realistic but also sufficiently diverse to simulate a wide range of scenarios.
@@ 5.3. Attention Black Hole
To address the challenge arising from the small proportion of the trigger image within the overall webpage screenshot, we produce Attention Black Hole (ABH), which optimizes the trigger image to effectively attract the model’s attention toward the trigger region.
@@
Minimizing this loss explicitly guides the model to focus on the trigger region, which we regard as the key to overcoming the challenge posed by the small size of the trigger image.
@@ 6. Experiments
First, as stated in the threat model, the primary goal of an attacker is to mislead the GUI agent into executing an incorrect action when exposed to the trigger image. We therefore design our first research question to evaluate the fundamental effectiveness of Chameleon in achieving this objective.
@@
Table 1. Datasets used in this work. Trigger Coverage Ratio denotes the percentage of the screenshot covered by the trigger image for each website.
Category Website URL Trigger Coverage Ratio
Shopping Amazon https://www.amazon.com/ 4.67%
Taobao https://www.taobao.com/ 7.59%
Social Media RedNote https://www.xiaohongshu.com/ 8.19%
Bilibili https://www.bilibili.com/ 9.97%
Music Streaming NetEase Cloud Music http://ir.music.163.com/en/ 8.17%
QQ Music https://y.qq.com/ 5.99%
Finally, we aim to understand the individual contributions of our core technical innovations. We therefore design our third research question to analyze the importance of each component within our framework.
@@ 6.1. Experimental Setup
Datasets. To ensure the realism of our evaluation, we selected six target websites spanning three representative categories of GUI-based online services, as summarized in Table 1. These categories reflect diverse user interaction patterns and GUI design paradigms, which are critical for assessing the robustness of Chameleon across practical use cases. For each website, we construct a validation set of 300 and a test set of 1,200 screenshot-instruction pairs using the LLM-Driven environment simulation introduced in Section 5.2. Importantly, the images and instructions in the training, validation, and test sets are mutually disjoint, ensuring rigorous evaluation. Figure 3 displays one representative screenshot for each of the six websites.
Figure 3. Representative screenshots for each website. Trigger images are outlined in red.
LVLMs for GUI agents. Our evaluation is conducted on a diverse set of four popular Large Vision-Language Models. These include two models specialized for GUI tasks, UI-TARS-7B-DPO (Qin et al., 2025) and OS-Atlas-Base-7B (Wu et al., 2024c), as well as two general-purpose LVLMs, Qwen2-VL-7B (Wang et al., 2024a) and LLaVA-1.5-13B (Liu et al., 2023). Notably, UI-TARS-7B-DPO and OS-Atlas-Base-7B are fine-tuned from Qwen2-VL-7B.
Baselines. We adopt a standard PGD-based adversarial attack as our baseline (Madry et al., 2017), representing methods utilized in previous research (Wu et al., 2024b; Aichberger et al., 2025). During the optimization of the trigger image, this baseline approach does not incorporate any environmental context, such as surrounding webpage content or stylistic information.
@@
goto [http://malicious-example.com]
@@
where $N_{\text{attack}}$ denotes the number of responses matching the target action within the test set, and $N_{\text{total}}$ represents the total size of the test set. A strict string matching criterion is adopted to verify whether the agent’s response matches the target action precisely. Consequently, our ASR metric provides a conservative estimate, representing a lower bound for the capability of the trigger image to pose practical security threats.
Implementation Details. We employ GPT-4o (Achiam et al., 2023) to parse the live website, preserving its core structural and stylistic components while removing existing content, and we use Qwen2.5-VL-32B-Instruct (Bai et al., 2025) to automatically generate realistic and diverse user instructions. The system prompt and user prompt for the GUI agent, as well as the agent’s action space, are adapted from VisualWebArena (Koh et al., 2024). Since it is challenging to collect authentic action histories, we follow ENVINJECTION (Wang et al., 2025) and randomly sample 0 to 10 historical actions for each instance. To ensure rigorous evaluation, we maintain a strict separation, with no overlap in the sampled action histories used across the training, validation, and testing. The hyperparameter $\lambda$ is set to $0.3$ and $\epsilon$ is set to $\frac{32}{255}$. The perturbation is optimized for a total of $5,000$ steps, with updates performed at each step using a fixed step size of $\alpha=\frac{1}{255}$. All experiments are conducted on a server equipped with 8 NVIDIA A100-PCIE-40GB GPUs.
@@ 6.2. RQ1: Effectiveness in Attacking Target GUI Agents
In this RQ, we evaluate whether Chameleon effectively misleads GUI agents into executing a predefined malicious action when exposed to an adversarial trigger image embedded within a target website.
Setup. We compare the ASR of Chameleon against the PGD baseline across four GUI agents on the six target websites detailed in our experimental setup.
Table 2. ASRs (%) of baseline and Chameleon across different GUI agents and websites. NetEase is used as an abbreviation for NetEase Cloud Music. The percentage shown at the top-right of each website name denotes the trigger coverage ratio (percentage of screenshot pixels occupied by the trigger image).
Model Method Shopping Social Media Music Streaming Avg.
Amazon4.67% Taobao7.59% RedNote8.19% Bilibili9.97% NetEase8.17% QQ Music5.99%
UI-TARS-7B-DPO Baseline 3.17 2.83 4.08 5.75 6.58 1.25 3.94
Chameleon 20.75 22.58 23.25 41.58 41.08 8.08 26.22
OS-Atlas-Base-7B Baseline 4.25 5.17 3.83 7.08 8.00 3.25 5.26
Chameleon 23.67 33.17 26.42 35.50 43.75 33.08 32.60
Qwen2-VL-7B Baseline 2.50 4.00 4.67 6.83 7.58 2.08 4.61
Chameleon 9.42 14.42 17.58 22.42 25.93 9.92 16.62
LLaVA-1.5-13B Baseline 4.75 6.25 7.08 8.17 8.75 3.50 6.42
Chameleon 37.17 43.00 60.75 50.83 76.58 32.08 50.07
Results. Table 2 presents the detailed ASR results for both our method and the baseline across all experimental configurations. Higher ASR indicates stronger attack effectiveness.
Analyses.Chameleon consistently outperforms the baseline by a substantial margin across all experimental configurations. While prior work (Wu et al., 2024b) has reported relatively high attack success rates for baseline methods when applied to static websites, Table 2 shows that the same baseline performs poorly in dynamically varying visual contexts, with ASRs remaining close to zero. In contrast, Chameleon achieves considerably higher ASRs across all settings, demonstrating its strong ability to reliably induce malicious behaviors in GUI agents under dynamic conditions. ❷ The effectiveness of Chameleon is positively correlated with the relative size of the trigger image within the screenshot. On websites such as NetEase Cloud Music, where the trigger image constitutes a larger portion of the visual area, the effectiveness is substantially higher, often exceeding 50%. Conversely, on websites such as Taobao, where the trigger image occupies a smaller region, the effectiveness remains below 20%. This finding is consistent with our earlier observation that when the trigger image represents only a small fraction of the overall screenshot, it becomes more difficult for the model to consistently attend to it.
Answer to RQ1: Chameleon demonstrates strong effectiveness in misleading GUI agents and consistently achieves substantially higher ASRs than baseline methods.

### v2

@@ 3. Dynamic-Environment Threat Model
Attacker’s Capabilities. Attackers can upload trigger images to a target website. Because website layouts and styles typically remain stable over short periods, attackers can reliably anticipate the overall structure and appearance of the target page. Furthermore, following prior work (Aichberger et al., 2025; Wu et al., 2024a), we assume that attackers have white-box access to the model powering the GUI agent  [^2] We also consider transfer-based black-box attacks in our evaluation., including knowledge of its gradients and architecture. We consider this assumption realistic because many GUI agents are deployed locally using open-weight LVLMs by users, allowing attackers to directly craft triggers against these publicly available models. Given the increasing deployment of GUI agents and the ease with which trigger images can be distributed online (Wang et al., 2024b; Tur et al., 2025), we think compromising even one popular model may lead to far-reaching consequences.
@@ 4. Methodology
Under the realistic dynamic-environment threat model, our results in Section 5.2 show that existing environmental injection attacks have very low effectiveness, and therefore fail to fully expose the vulnerabilities of GUI agents in real-world settings.
To tackle the challenge posed by dynamic environments, we propose Chameleon, a novel attack framework for LVLM-powered GUI agents. We begin by describing the overall attack pipeline in Section 4.1. We then present LLM-Driven Environment Simulation in Section 4.2, which automatically generates large-scale realistic training data that captures dynamic webpage variations. Finally, we introduce Attention Black Hole in Section 4.3, which ensure more effective training by encouraging the model to consistently focus its attention on the trigger region. An overview of Chameleon is shown in Figure 3.
@@ 4.1. Overview
For a chosen target website, such as a social media platform like RedNote  [^3] https://www.xiaohongshu.com/ or an e-commerce site like Amazon  [^4] https://www.amazon.com/, the attacker first selects a benign-looking trigger image $I$, such as a product photo or a post cover. This image serves as the starting point and will be optimized during training.
At each training step, we first use our LLM-Driven Environment Simulation $G_{\mathrm{LES}}$ (Section 4.2) to construct training data. Specifically, it generates a webpage screenshot $s$ that contains the perturbed trigger image, a corresponding user instruction $p_{u}$, and a pixel-level mask $mask^{shot}$ that identifies the trigger region within $s$. We then update the perturbation $\delta$ inside the trigger region using Projected Gradient Descent (PGD) (Madry et al., 2017). To ensure the resulting malicious trigger image remains imperceptible to users, we constrain the perturbation $\delta$ to a predefined $\ell_{\infty}$ norm bound $\epsilon$, formally:
@@
where $\lambda$ controls the relative weight of the attention loss.
After the training procedure described above, the attacker uploads the resulting trigger image to the target website, acting as a regular user. Once the trigger image is integrated into a public post or product listing, any GUI agent browsing the site and encountering the trigger image within a webpage will be susceptible to attack, potentially performing the intended harmful action. Algorithm 1 summarizes the overall procedure of Chameleon, covering the training and deployment stages.
@@
$C\leftarrow$ Crawl large-scale multimodal content;
@@ 4.2. LLM-Driven Environment Simulation
We argue that enabling the trigger to generalize across dynamic environments requires training over a large collection of realistic visual environments. To address this challenge, we introduce LLM-Driven Environment Simulation (LES). LES leverages the generative capabilities of LLMs to automatically synthesize a large number of realistic training samples. By systematically varying the trigger’s position and its surrounding visual context during synthesis, LES ensures the optimized trigger learns to remain effective under continuously changing environments. Concretely, we define a generation function $G_{LES}$ that outputs a rendered webpage screenshot, a binary mask indicating the trigger location, and a user instruction for downstream training.
First, to ensure that our simulated environments are realistic, we construct a high-fidelity HTML template $T$ for each target website. We begin by collecting the raw HTML from the live website and treat it as the initial reference for both layout and styling. Our goal is to remove concrete content (e.g., product titles, prices, user posts, and recommendations) while preserving the core structural and stylistic components (e.g., DOM hierarchy, CSS classes, layout containers, and visual themes), so that the resulting template remains visually indistinguishable from the real site after being populated. To achieve this, we leverage a flagship LLM in a multi-turn, interactive refinement process. In each round, we provide the LLM with the raw HTML and instruct it to abstract away concrete content into clearly defined placeholders that can later be filled with arbitrary image-text pairs. We then render the produced template, inspect whether any real content is still hard-coded or whether the layout has been unintentionally altered, and feed this feedback back to the LLM for further refinement. After several rounds, this process yields a clean, fillable HTML template $T$ that preserves the authentic look of the target website while allowing us to programmatically inject diverse content during simulation. Notably, for each target website, this template construction is a one-time effort, and a single HTML template $T$ can be reused to generate all simulated environments for that website.
We then construct a large-scale, multimodal corpus $C$ by crawling the target website for actual content, gathering an average of over 5,000 distinct items (e.g., product images, titles, prices) for each website. This process ensures our content pool is not only realistic but also sufficiently diverse to simulate a wide range of scenarios.
@@ 4.3. Attention Black Hole
After obtaining large-scale realistic training data, the next challenge is how to optimize the trigger more effectively. We argue that reliable attacks in dynamically changing visual environment require the model to consistently prioritize the trigger over competing interface elements. Thus, we propose Attention Black Hole (ABH). By converting attention weights into explicit supervisory signals, ABH encourages the agent to consistently focus on the trigger region, thereby improving robustness by reducing interference from distracting content in the dynamic environment.
@@
Minimizing this loss explicitly guides the model to focus on the trigger region, which is crucial for disregarding the distractions caused by the dynamic environment and maintaining attack robustness.
@@ 5. Experiments
Figure 4. Illustrative screenshots for each website. Trigger images are outlined in red.
First, as stated in our proposed threat model, the primary goal of an attacker is to mislead the GUI agent into executing an incorrect action when exposed to the trigger image. We therefore design our first research question to evaluate the fundamental effectiveness of Chameleon in achieving this objective.
@@
Thirdly, we aim to understand the individual contributions of our core technical innovations. We therefore design our third research question to analyze the importance of each component within our framework.
@@
Finally, we explore whether existing defenses can mitigate Chameleon under our realistic dynamic-environment threat model, and what practical mitigation strategies may be effective. We therefore design our fourth research question to study potential defenses.
RQ4: What defenses are effective against Chameleon under the dynamic-environment threat model? We evaluate a set of practical defense strategies and measure their impact on attack success rate, aiming to identify promising directions for securing GUI agents in realistic deployments.
@@ 5.1. Experimental Setup
Datasets. To ensure the realism of our evaluation, we selected six target websites spanning three representative categories of GUI-based online services, as summarized in Table 1. These categories reflect diverse user interaction patterns, which are critical for assessing the effectiveness of Chameleon across practical use cases. For each website, we construct a validation set of 300 and a test set of 1,200 screenshot-instruction pairs using the LLM-Driven Environment Simulation introduced in Section 4.2. Importantly, the images and instructions in the training, validation, and test sets are mutually disjoint, ensuring rigorous evaluation. Figure 4 displays some example screenshots for each of the six websites.
Table 1. Datasets used in this work.
Category Website URL
Shopping Amazon https://www.amazon.com/
Taobao https://www.taobao.com/
Social Media RedNote https://www.xiaohongshu.com/
Bilibili https://www.bilibili.com/
Music Streaming NetEase Cloud Music http://ir.music.163.com/en/
QQ Music https://y.qq.com/
LVLMs for GUI agents. Our evaluation is conducted on a diverse set of four popular and representative Large Vision-Language Models. These include two models specialized for GUI tasks, UI-TARS-7B-DPO (Qin et al., 2025) and OS-Atlas-Base-7B (Wu et al., 2024b), as well as two general-purpose LVLMs, Qwen2-VL-7B (Wang et al., 2024a) and LLaVA-1.5-13B (Liu et al., 2023). Notably, UI-TARS-7B-DPO and OS-Atlas-Base-7B are fine-tuned from Qwen2-VL-7B.
Baselines. Although a large body of prior work (Lu et al., 2025; Zhao et al., 2025; Wang et al., 2025; Aichberger et al., 2025) studies environmental injection attacks, we find that most of them cannot be directly applied under our dynamic-environment threat model. Many approaches assume an attacker with administrative privileges that are not practical, while others rely on explicit textual injections that are easily filtered. After carefully filtering the literature to match our attacker capabilities and deployment constraints, we select two representative baselines:
  • PGD (Wu et al., 2024a). We adopt a standard PGD-based adversarial attack (Madry et al., 2017) as a baseline, following prior trigger-image EIA studies such as (Wu et al., 2024a). During trigger optimization, this baseline does not incorporate any environmental context, including surrounding webpage content or stylistic information.
  • MIP (Aichberger et al., 2025). We implement MIP, which constructs training samples by manually creating dynamic environments and then optimizes the trigger using PGD. Following its experimental setup, for each trigger image we manually craft 12 screenshots with varying contexts for optimization.
@@
goto [http://one-example.com]
@@
where $N_{\text{attack}}$ denotes the number of responses matching the target action within the test set, and $N_{\text{total}}$ represents the total size of the test set. A strict string matching criterion is adopted to verify whether the agent’s response matches the target action precisely.
Implementation Details. We employ GPT-4o (Achiam et al., 2023) to parse the live website, preserving its core structural and stylistic components while removing existing content, and we use Qwen2.5-VL-32B-Instruct (Bai et al., 2025) to automatically generate realistic and diverse user instructions. The system prompt and user prompt for the GUI agent, as well as the agent’s action space, are adapted from VisualWebArena (Koh et al., 2024). Since it is challenging to collect authentic action histories, we follow ENVINJECTION (Wang et al., 2025) and randomly sample 0 to 10 historical actions for each instance. To ensure rigorous evaluation, we maintain a strict separation, with no overlap in the sampled action histories used across the training, validation, and testing. The hyperparameter $\lambda$ is set to $0.3$ and $\epsilon$ is set to $\frac{32}{255}$. The perturbation is optimized for a total of $5,000$ steps, with updates performed at each step using a fixed step size of $\alpha=\frac{1}{255}$.
Table 2. ASRs (%) of baselines and Chameleon across different GUI agents and websites. NetEase is used as an abbreviation for NetEase Cloud Music. Gain denotes the improvement of Chameleon over the best-performing baseline for each setting.
Model Method Shopping Social Media Music Streaming Avg.
Amazon Taobao RedNote Bilibili NetEase QQ Music
UI-TARS-7B-DPO PGD 3.17 2.83 4.08 5.75 6.58 1.25 3.94
MIP 5.67 4.42 5.17 7.17 11.08 1.33 5.78
Chameleon 20.75 22.58 23.25 41.58 41.08 8.08 26.22
Gain $\times$3.66 $\times$5.11 $\times$4.50 $\times$5.80 $\times$3.71 $\times$6.08 $\times$4.54
OS-Atlas-Base-7B PGD 4.25 5.17 3.83 7.08 8.00 3.25 5.26
MIP 6.17 7.25 3.92 9.92 10.08 3.50 6.81
Chameleon 23.67 33.17 26.42 35.50 43.75 33.08 32.60
Gain $\times$3.84 $\times$4.58 $\times$6.74 $\times$3.58 $\times$4.34 $\times$9.45 $\times$4.79
Qwen2-VL-7B PGD 2.50 4.00 4.67 6.83 7.58 2.08 4.61
MIP 6.75 8.17 7.08 11.50 14.17 4.17 8.64
Chameleon 9.42 14.42 17.58 22.42 25.93 9.92 16.62
Gain $\times$1.40 $\times$1.76 $\times$2.48 $\times$1.95 $\times$1.83 $\times$2.38 $\times$1.92
LLaVA-1.5-13B PGD 4.75 6.25 7.08 8.17 8.75 3.50 6.42
MIP 9.17 10.17 7.92 14.67 15.25 9.42 11.1
Chameleon 37.17 43.00 60.75 50.83 76.58 32.08 50.07
Gain $\times$4.05 $\times$4.23 $\times$7.67 $\times$3.46 $\times$5.02 $\times$3.41 $\times$4.51
@@ 5.2. RQ1: Effectiveness in Attacking Target GUI Agents
In this RQ, we evaluate whether Chameleon can reliably mislead GUI agents into executing a predefined malicious action when an adversarial trigger is embedded in a target website under the dynamic-environment threat model.
Setup. We compare the ASR of Chameleon with two baselines across four representative GUI agents and six target websites described in our experimental setup.
Results. Table 2 reports the ASRs (%) of all methods across models and websites. Higher ASR indicates stronger attack effectiveness.
Analyses.Existing attacks fail to fully expose GUI agent vulnerabilities under realistic dynamic environments. Although prior studies reported relatively high ASRs for baseline attacks under static or slightly varying settings (Wu et al., 2024a; Aichberger et al., 2025), Table 2 shows that both PGD and MIP remain largely ineffective in our dynamic-environment evaluation, with ASRs remaining close to zero across models and websites. For example, on OS-Atlas-Base-7B, PGD achieves only 5.26% ASR on average and MIP reaches 6.81%, suggesting that these approaches substantially underestimate the vulnerability of GUI agents when evaluated in realistic, continuously changing web contexts. ❷ Chameleon achieves substantially higher attack effectiveness across all configurations. In contrast to the baselines, Chameleon consistently improves ASR by a large margin for every evaluated agent and website. For instance, on OS-Atlas-Base-7B, Chameleon increases the average ASR from 5.26% (PGD) and 6.81% (MIP) to 32.60%. Similar gains are observed on UI-TARS-7B-DPO (from 3.94% and 5.78% to 26.22%) and Qwen2-VL-7B (from 4.61% and 8.64% to 16.62%). These results demonstrate that Chameleon can effectively uncover previously hidden vulnerabilities of GUI agents that remain unexposed by existing attacks under realistic dynamic environments.
Answer to RQ1: Under the dynamic-environment threat model, existing attacks exhibit very low effectiveness and thus fail to fully reveal GUI agent vulnerabilities, whereas Chameleon achieves substantially higher ASRs across all models and websites.